The average of three numbers is 77 . The
first number and second number is 78 and 79 respectively. Find the third number

Answer by Cromlix(4381)

total of 3 numbers = 3 * 77 = 231
Total of known numbers = 78 + 79 = 157
231 - 157 = 74
Therefore 74 is the third number.
Hope this helps.
:-)